# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v2

@@ 2.1 Model and ex–ante objective in estimated diffusion markets
2.3 Gaussian references at a fixed decision time
This subsection collects Gaussian benchmarks that yield closed-form decision-time reference allocations. Fix a decision time $t\in[0,T)$ and suppose an external procedure provides an $\mathcal{F}_{t}$-measurable input law $q_{t}$ for the uncertain premium parameter/state. We treat $q_{t}$ as a decision-time input (not a belief process updated within a run): it is held fixed over the remaining horizon $[t,T]$ and the resulting references are used only as validation targets and controlled diagnostics (not as characterizations of the unrestricted optimum). If the decision is revisited at a later time $t^{\prime}$, the reference can be recomputed by resolving the decision-time problem with the new input $q_{t^{\prime}}$ (plug-in replanning).
We present two decision-time references. Section 2.3.1 considers static drift uncertainty and provides the controlled benchmark used in the high-dimensional scaling/geometry experiments of Section 4. Section 2.3.2 considers a mean-reverting (OU) premium factor with Gaussian decision-time uncertainty for the current factor level, which induces a horizon-dependent effective Gaussian law for the time-averaged premium and yields a tractable reference used in the hedging-demand recovery study of Section 5. Appendix A gives a linear–Gaussian illustration of how an external filtering routine may produce a time-indexed input law $q_{t}$; it is included only as an input mechanism (plug-in replanning), not as a belief-state control formulation.
@@ 2.3.1 Static Gaussian drift uncertainty
We start from a time-homogeneous Gaussian benchmark in which the (vector) risk premium is an unobserved static parameter. This “unknown mean” specification is a standard device for isolating estimation risk/parameter uncertainty in portfolio choice (e.g., Kandel and Stambaugh, 1996; Barberis, 2000; Pástor, 2000). Fix a decision time $t\in[0,T)$ and remaining horizon $\tau:=T-t$. The agent chooses a $\theta$-blind policy on $[t,T]$, and all decision-time uncertainty about the (time-constant) premium is summarized by an $\mathcal{F}_{t}$-measurable law $q_{t}(d\theta)$, which we do not update over $[t,T]$.
Market model (static latent drift).
@@
where $\Sigma\in\mathbb{R}^{d\times d}$ is symmetric positive definite and $\theta\in\mathbb{R}^{d}$ is a latent constant excess-return vector. At decision time $t$, uncertainty about $\theta$ is represented by the input law
@@ A $\theta$-blind portfolio fraction process $\pi_{s}\in\mathbb{R}^{d}$ generates wealth on $[t,T]$:
We evaluate the (conditional) decision-time objective by mixing over $q_{t}$:
$J_{t}(\pi):=\mathbb{E}_{\theta\sim q_{t}}\Big[\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta,\,\mathcal{F}_{t}\big]\Big].$ (24)
Under full information and CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ ($\gamma>0,\gamma\neq 1$), the oracle Merton rule is $\pi^{\star}(\theta)=\frac{1}{\gamma}\Sigma^{-1}\theta$ (Merton, 1969, 1971), which is infeasible here because $\theta$ is latent.
Analytic $q_{t}$-references via portfolios constant over $[t,T]$. To obtain transparent closed-form targets, we temporarily restrict attention to portfolio fractions that are constant in trading time $s$ on $[t,T]$:
$\pi_{s}\equiv\pi\in\mathbb{R}^{d},\hskip 18.49988pts\in[t,T].$ (25)
The superscript $\mathrm{const}$” used below refers to this constancy in $s$. The resulting reference may still depend on the decision time $t$ through the input $q_{t}$ and on the remaining horizon $\tau=T-t$. (Decision-time constant-portfolio targets of this kind are closely connected to myopic/plug-in benchmarks used in the parameter-uncertainty literature; see, e.g., Barberis (2000); Pástor (2000).)
CRRA ($\gamma>1$): tilted optimality and Gaussian shrinkage.
For $\gamma>1$ and constant $\pi$, conditional on $\theta$ the terminal wealth is lognormal and
where $M_{q_{t}}(u):=\mathbb{E}_{\theta\sim q_{t}}[\exp(u^{\top}\theta)]$ is the moment generating function of $q_{t}$. Any interior optimizer $\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)$ satisfies the tilted first-order condition
$\gamma\,\Sigma\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\nabla_{u}\log M_{q_{t}}(u)\Big|_{u=(1-\gamma)\tau\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)}.$ (27)
@@
then $\nabla_{u}\log M_{q_{t}}(u)=m_{t}+P_{t}u$ and the reference reduces to the linear system
$\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=m_{t},$ (29)
hence
$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)^{-1}m_{t}.$ (30)
For $\gamma>1$, this takes the familiar shrinkage form
$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{t}\big)^{-1}m_{t},\hskip 18.49988pt(\gamma>1),$ (31)
which makes explicit how decision-time uncertainty $P_{t}$ dampens exposures as the remaining horizon grows (a qualitative effect widely emphasized in return-predictability/parameter-uncertainty studies; see, e.g., Barberis (2000); Pástor (2000); Kandel and Stambaugh (1996)).
@@ 2.3.2 Mean-reverting Gaussian premium and an induced horizon-dependent reference
We next replace the static premium by a mean-reverting Gaussian premium factor, a standard reduced-form device for return predictability and intertemporal hedging (Campbell and Viceira, 2002; Xia, 2001). Fix a decision time $t\in[0,T)$ and remaining horizon $\tau:=T-t$. At time $t$ we take as input a Gaussian law for the current factor level $Y_{t}$ (rather than a point estimate), capturing decision-time statistical uncertainty (e.g., interval-type uncertainty quantification, resampling variability, or asymptotic approximations). Holding this input fixed over $[t,T]$, OU propagation and time-averaging induce a horizon-dependent effective premium law; when $\rho\neq 0$, return–factor shock correlation adds a further quadratic term relevant for hedging demand (Campbell and Viceira, 2002; Xia, 2001).
OU premium dynamics.
Let $Y_{s}\in\mathbb{R}^{m}$ be a premium factor following
@@
where $\bar{y}\in\mathbb{R}^{m}$ and $\Xi\in\mathbb{R}^{m\times m}$. Risky excess returns satisfy
@@
Decision-time input law for the factor.
At decision time $t$, we take as input an $\mathcal{F}_{t}$-measurable Gaussian law for the current factor level,
@@
A common special case is purely model-implied predictive propagation from an initial uncertainty $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$ without conditioning on observations:
$m_{t}=\bar{y}+e^{-Kt}(m_{0}-\bar{y}),$
$P_{t}=e^{-Kt}P_{0}e^{-K^{\top}t}+\int_{0}^{t}e^{-K(t-s)}\,\Xi\Xi^{\top}\,e^{-K^{\top}(t-s)}\,ds,$ (36)
which is an OU transition (prior-predictive) law and should not be interpreted as an information update. (References for return predictability and OU-style investment-opportunity dynamics include Campbell and Viceira (2002); Xia (2001).)
Induced Gaussian law for the time-averaged premium over $[t,T]$.
Define the remaining-horizon integrated factor and the associated horizon-averaged effective premium:
$I_{t,T}:=\int_{t}^{T}Y_{s}\,ds\in\mathbb{R}^{m},\hskip 18.49988pt\bar{\theta}_{t,\tau}:=\frac{1}{\tau}\,B\,I_{t,T}\in\mathbb{R}^{d}.$ (37)
Let
$A(\tau):=K^{-1}\big(I-e^{-K\tau}\big)\in\mathbb{R}^{m\times m}.$
Under the linear–Gaussian OU model, $I_{t,T}$ is Gaussian under the input law (35):
$I_{t,T}\sim\mathcal{N}\big(m_{I}(t,\tau),C_{I}(t,\tau)\big),$
with
$m_{I}(t,\tau)=\mathbb{E}[I_{t,T}]=\tau\bar{y}+A(\tau)\,(m_{t}-\bar{y}),$ (38)
$C_{I}(t,\tau):=\mathrm{Cov}(I_{t,T})=A(\tau)\,P_{t}\,A(\tau)^{\top}+\int_{0}^{\tau}A(u)\,\Xi\Xi^{\top}\,A(u)^{\top}\,du.$ (39)
Consequently, the horizon-averaged effective premium is Gaussian,
$\bar{\theta}_{t,\tau}\sim\mathcal{N}\!\big(m_{\bar{\theta}}(t,\tau),P_{\bar{\theta}}(t,\tau)\big),\hskip 18.49988ptm_{\bar{\theta}}(t,\tau)=\frac{1}{\tau}B\,m_{I}(t,\tau),\hskip 18.49988ptP_{\bar{\theta}}(t,\tau)=\frac{1}{\tau^{2}}B\,C_{I}(t,\tau)\,B^{\top}.$ (40)
The pair $\big(m_{\bar{\theta}}(t,\tau),P_{\bar{\theta}}(t,\tau)\big)$ is the horizon-dependent effective Gaussian law that we use as a controlled analytic input.
Closed-form decision-time references under portfolios constant over $[t,T]$.
As in Section 2.3.1, we define an analytic reference by restricting to portfolio fractions that are constant in trading time $s$ over $[t,T]$, i.e. $\pi_{s}\equiv\pi\in\mathbb{R}^{d}$. The superscript $\mathrm{const}$” below refers to this constancy in $s$; the resulting reference may still depend on the decision time $t$ through $(m_{t},P_{t})$ and on the remaining horizon $\tau$ through OU averaging.
CRRA ($\gamma>1$).
When $\rho\neq 0$, return shocks and factor shocks are contemporaneously correlated, producing an additional quadratic term in the constant-portfolio objective (hedging-demand channel; see, e.g., Campbell and Viceira (2002); Xia (2001)). Define the cross-covariance between $I_{t,T}$ and the return Brownian increment $W_{T}-W_{t}$:
$C_{IW}(\tau):=\mathrm{Cov}\!\big(I_{t,T},\,W_{T}-W_{t}\big)=\int_{0}^{\tau}A(u)\,\Xi\,\rho^{\top}\,du\;\in\;\mathbb{R}^{m\times d},$ (41)
and the induced symmetric cross term
Then the Gaussian decision-time reference under constant portfolios is characterized by
$\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=B\,m_{I}(t,\tau),\hskip 18.49988pt(\gamma>1).$ (43)
Equivalently,
$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)^{-1}B\,m_{I}(t,\tau).$
When $\rho=0$, we have $C_{IW}(\tau)=0$ and $M_{\mathrm{cross}}(\tau)=0$, and the reference reduces to the independence-case shrinkage form obtained by plugging the effective Gaussian law (40) into the static Gaussian formula of Section 2.3.1:
$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{\bar{\theta}}(t,\tau)\big)^{-1}m_{\bar{\theta}}(t,\tau),\hskip 18.49988pt(\rho=0,\ \gamma>1).$
@@ 3 Pontryagin–Guided Policy Optimization under Latent Parameter Uncertainty
Our solver has two components. Stage 1 (PG–DPO) performs stochastic gradient ascent on $J(\varphi)=\mathbb{E}[U(X_{T}^{\pi_{\varphi},\theta})]$ by differentiating through an Euler simulator via BPTT. Stage 2 (P–PGDPO) aggregates the BPTT costate blocks across $\theta\sim q$ and enforces the $q$-aggregated Pontryagin first-order condition, yielding a deployable projected control.
Section 3.1 establishes the conditional BPTT–PMP correspondence and specifies the adjoint blocks needed for projection. Section 3.2 develops the $q$-aggregated projection and a residual-based policy-gap bound. Practical coupling mechanisms (residual-form projection and interactive distillation) are deferred to Appendix F.
3.1 PG–DPO as stochastic gradient ascent and a conditional BPTT–PMP correspondence
@@
$\pi_{\varphi}:\ [0,T]\times(0,\infty)\times\mathbb{R}^{m}\to\mathbb{R}^{d},\hskip 18.49988pt(t,x,y)\mapsto\pi_{\varphi}(t,x,y),\hskip 18.49988pt\varphi\in\mathbb{R}^{p},$ (44)
which does not take $\theta$ as an input. Conditional on a fixed frozen $\theta$, the controlled state $(X_{t}^{\pi,\theta},Y_{t}^{\theta})_{t\in[0,T]}$ evolves as
$\frac{dX_{t}^{\pi,\theta}}{X_{t}^{\pi,\theta}}$ $=\Big(r+\pi_{t}^{\top}b\big(Y_{t}^{\theta},\theta\big)\Big)\,dt+\pi_{t}^{\top}\sigma\big(Y_{t}^{\theta},\theta\big)\,dW_{t},\hskip 18.49988ptX_{0}=x>0,$ (45)
@@
Discretization, sampling over $\theta$, and baseline PG–DPO update.
@@
$\big(x^{(i)},y^{(i)},\theta^{(i)},\{\Delta W_{k}^{(i)},\Delta W_{k}^{Y,(i)}\}_{k=0}^{N-1},\varphi\big)\longmapsto U\big(X_{N}^{(i)}\big)$
@@
A typical PG–DPO update samples a mini-batch $\{z_{0}^{(i)}\}_{i=1}^{M}\sim\nu$, samples latent parameters either independently per episode ($\theta^{(i)}\sim q$) or shared within the batch (one $\theta\sim q$ reused across $i$), simulates Euler rollouts, and performs stochastic gradient ascent. Suppressing episode indices and writing $\theta$ for the frozen draw used in the rollout, the Euler recursion is
@@
starting from $(t_{0},X_{0},Y_{0})=(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})$. The episode reward is
@@
$G_{t_{k}}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t_{k}\}\big)\qquad\text{(with the usual augmentation)},$
@@
$p_{k}^{\Delta t,\theta}:=\mathbb{E}[p_{k}\mid G_{t_{k}}^{\theta}],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}[p_{x,k}\mid G_{t_{k}}^{\theta}],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}[p_{y,k}\mid G_{t_{k}}^{\theta}].$

### v3

@@ 2.1 Model and ex–ante objective in estimated diffusion markets
2.3 Gaussian decision-time references
This subsection collects Gaussian benchmarks that yield closed-form decision-time reference allocations. Fix a decision time $t\in[0,T)$ and let an external routine provide an $\mathcal{F}_{t}^{\mathrm{obs}}$-measurable input law $q_{t}$ for the uncertain premium parameter/state. We treat $q_{t}$ as a decision-time input (not a belief process controlled/updated within a run): it is held fixed over the remaining horizon $[t,T]$. The resulting portfolios are used only as validation targets and controlled diagnostics (not as characterizations of the unrestricted optimum). If the decision is revisited at a later time $t^{\prime}$, the reference is recomputed by resolving the decision-time problem with the new input $q_{t^{\prime}}$ (plug-in replanning).
We present two references. Section 2.3.1 considers static drift uncertainty and provides the benchmark used in the high-dimensional scaling/geometry experiments of Section 4. Section 2.3.2 considers a mean-reverting (OU) premium factor with Gaussian decision-time uncertainty for the current factor level; OU propagation and time-averaging induce a horizon-dependent effective Gaussian premium law, yielding a tractable reference used in the hedging-demand recovery study of Section 5. All derivations are deferred to Appendix A.
@@ 2.3.1 Static Gaussian drift uncertainty
Benchmark.
@@
where $\Sigma\in\mathbb{R}^{d\times d}$ is symmetric positive definite and $\theta\in\mathbb{R}^{d}$ is a latent constant excess-return vector. At decision time $t$, uncertainty about $\theta$ is summarized by the input law
@@ A $\theta$-blind portfolio fraction process $\pi_{s}\in\mathbb{R}^{d}$ generates wealth on $[t,T]$:
We evaluate the decision-time objective by mixing over $q_{t}$:
$J_{t}(\pi):=\mathbb{E}_{\theta\sim q_{t}}\Big[\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta,\,\mathcal{F}_{t}^{\mathrm{obs}}\big]\Big].$ (27)
Decision-time constant-portfolio reference.
To obtain a transparent closed-form target, we restrict to portfolio fractions constant over $[t,T]$:
$\pi_{s}\equiv\pi\in\mathbb{R}^{d},\hskip 18.49988pts\in[t,T],$ (28)
and take CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$ with $\gamma>1$.
@@
then the constant-portfolio reference is
$\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{t}\big)^{-1}m_{t},\hskip 18.49988pt\tau:=T-t,\qquad(\gamma>1).$ (30)
This exhibits the familiar Gaussian shrinkage: decision-time uncertainty $P_{t}$ dampens exposures, increasingly so as the remaining horizon $\tau$ grows.
@@ 2.3.2 Mean-reverting Gaussian premium and an induced horizon-dependent reference
OU premium factor.
Let $Y_{s}\in\mathbb{R}^{m}$ follow
@@
and let risky excess returns satisfy
@@
Decision-time input law and induced effective premium law.
At decision time $t$, we take as input a Gaussian law for the current factor level,
@@
Define the remaining-horizon integrated factor and horizon-averaged effective premium:
$I_{t,T}:=\int_{t}^{T}Y_{s}\,ds\in\mathbb{R}^{m},\hskip 18.49988pt\bar{\theta}_{t,\tau}:=\frac{1}{\tau}\,B\,I_{t,T}\in\mathbb{R}^{d},\hskip 18.49988pt\tau:=T-t.$ (35)
Under the linear–Gaussian OU model, holding the input law (34) fixed over $[t,T]$ induces a horizon-dependent Gaussian law
$\bar{\theta}_{t,\tau}\sim\mathcal{N}\!\big(m_{\bar{\theta}}(t,\tau),P_{\bar{\theta}}(t,\tau)\big),$ (36)
where explicit formulas for $m_{\bar{\theta}}(t,\tau)$ and $P_{\bar{\theta}}(t,\tau)$ are given in Appendix A.
Decision-time constant-portfolio reference (CRRA, $\gamma>1$).
Restricting to constant portfolio fractions $\pi_{s}\equiv\pi$ on $[t,T]$ as above, the Gaussian decision-time reference under CRRA ($\gamma>1$) is characterized by the linear system
$\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)\,\pi_{q_{t},\gamma}^{\mathrm{const}}(t,\tau)=B\,m_{I}(t,\tau),\hskip 18.49988pt(\gamma>1),$ (37)
where $m_{I}(t,\tau):=\mathbb{E}[I_{t,T}]$, $C_{I}(t,\tau):=\mathrm{Cov}(I_{t,T})$, and the cross term $M_{\mathrm{cross}}(\tau)$ captures the effect of instantaneous return–factor correlation $\rho\neq 0$ (hedging-demand channel). Closed-form expressions for $m_{I}(t,\tau)$, $C_{I}(t,\tau)$, and $M_{\mathrm{cross}}(\tau)$ are collected in Appendix A.
When $\rho=0$, we have $M_{\mathrm{cross}}(\tau)=0$ and the reference reduces to the independence-case shrinkage form obtained by plugging the effective Gaussian law (36) into the static Gaussian formula (30).
@@ 3 Pontryagin–Guided Policy Optimization under Latent Parameter Uncertainty
We use a deterministic policy network $\pi_{\varphi}:\mathcal{D}\to\mathbb{R}^{d}$, parameterized by $\varphi$, where $\mathcal{D}\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$ denotes a working domain inside the state space.
Algorithmic overview and terminology. Our full solver is Pontryagin–guided direct policy optimization (PG–DPO) and consists of two stages. Stage 1 (DPO) performs stochastic gradient ascent on the ex–ante objective $J(\varphi)=\mathbb{E}\big[U\big(X_{T}^{\pi_{\varphi},\theta}\big)\big]$ by differentiating through an Euler simulator via BPTT. Stage 2 (Pontryagin guidance / projection) aggregates the BPTT adjoint (costate) blocks across $\theta\sim q$ and enforces the $q$-aggregated Pontryagin first-order condition, yielding a deployable projected control. We refer to Stage 1 alone as DPO and to the two-stage procedure (Stage 1+2) as PG–DPO.
Section 3.1 establishes the conditional BPTT–PMP correspondence and specifies the adjoint blocks needed for the projection stage. Section 3.2 develops the $q$-aggregated projection and a residual-based policy-gap bound. Practical coupling mechanisms (residual-form projection and interactive distillation) are deferred to Appendix G.
3.1 DPO via stochastic gradient ascent and a conditional BPTT–PMP correspondence
@@
$\pi_{\varphi}:\ \mathcal{D}\to\mathbb{R}^{d},\hskip 18.49988pt\mathcal{D}\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m},\hskip 18.49988pt(t,x,y)\mapsto\pi_{\varphi}(t,x,y),\hskip 18.49988pt\varphi\in\mathbb{R}^{p},$ (38)
which does not take $\theta$ as an input. Conditional on a fixed frozen $\theta$, the controlled state $(X_{t}^{\pi_{\varphi},\theta},Y_{t}^{\theta})_{t\in[0,T]}$ evolves as
$\frac{dX_{t}^{\pi_{\varphi},\theta}}{X_{t}^{\pi_{\varphi},\theta}}$ $=\Big(r+\pi_{t}^{\top}b\big(Y_{t}^{\theta},\theta\big)\Big)\,dt+\pi_{t}^{\top}\sigma\big(Y_{t}^{\theta},\theta\big)\,dW_{t},\hskip 18.49988ptX_{0}=x>0,$ (39)
@@
Discretization, sampling over $\theta$, and the baseline DPO update.
@@
$\big(x_{0}^{(i)},y_{0}^{(i)},\theta^{(i)},\{\Delta W_{k}^{(i)},\Delta W_{k}^{Y,(i)}\}_{k=0}^{N-1},\varphi\big)\longmapsto U\big(X_{N}^{(i)}\big)$
@@
A typical DPO update samples a mini-batch $\{z_{0}^{(i)}\}_{i=1}^{M}\sim\nu$, samples latent parameters either independently per episode ($\theta^{(i)}\sim q$) or shared within the batch (one $\theta\sim q$ reused across $i$), simulates Euler rollouts, and performs stochastic gradient ascent. Suppressing episode indices and writing $\theta$ for the frozen draw used in the rollout, the Euler recursion is
@@
starting from $(t_{0},X_{0},Y_{0})=(t_{0}^{(i)},x_{0}^{(i)},y_{0}^{(i)})$.
Remark 4 (Positivity of wealth in the discrete simulator).
In continuous time, the wealth process in (39) stays strictly positive for any admissible portfolio. A naive additive Euler step may nevertheless generate negative $X_{k}$ for coarse steps or aggressive controls, making CRRA utilities (and the higher-order derivatives used in BPTT) ill-defined. In implementations we therefore use a positivity-preserving discretization (e.g. exponential Euler on log-wealth) or apply a mild truncation/clipping so that $X_{k}\geq\underline{x}>0$ on the computation graph. The correspondence results below should be read for such a positivity-preserving simulator; equivalently, they apply on the event $\{\min_{0\leq k\leq N}X_{k}>0\}$.
The episode reward is
@@
$\mathcal{G}_{t_{k}}^{\theta}:=\sigma\!\big(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t_{k}\}\big)\qquad\text{(with the usual augmentation)},$
@@
$p_{k}^{\Delta t,\theta}:=\mathbb{E}[p_{k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}[p_{x,k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}[p_{y,k}\mid\mathcal{G}_{t_{k}}^{\theta}].$